OpenStudy (anonymous):

Identify all values of t E (0,2pi] that satisfy csc(t)= 2 I need an example to help me solve this … please

OpenStudy (anonymous):

\[\csc(t)=\frac{1}{\sin(t)}\] \[\csc(t)=2\iff \frac{1}{\sin(t)}=2\iff \sin(t)=\frac{1}{2}\] is a start